OpenStudy (anonymous):

Simplify completely 25w^6 over 10w^3 divided by 30w^2 over 5w?

OpenStudy (lgbasallote):

\[\huge \frac{\frac{25w^6}{10w^3}}{\frac{30w^2}{5w}} \implies \frac{\frac{5w^3}{2}}{6w}\] does that help?

OpenStudy (anonymous):

so which one is your final answer?

OpenStudy (anonymous):

\[25w ^{6/10}*w ^{3}/30w ^{2/5}*w\]

OpenStudy (anonymous):

is this the question

OpenStudy (lgbasallote):

i gave you no answer...i gave you a few steps to start you off

OpenStudy (anonymous):

yea thats the question

OpenStudy (anonymous):

and oh thanks:)

OpenStudy (lgbasallote):

it's illegal here to give answers...i wouldnt want to get suspended again:/

OpenStudy (anonymous):